Ag2co3 is converted to Ag20 + CO2 is the forward reaction endothermic?
1
Expert's answer
2016-02-12T05:37:41-0500
Ag2CO3 = Ag2O + CO2
dH = (dH(Ag2CO3) + dH(CO2))-dH(Ag2CO3)=(-31.1 + -393,51) - (-506,1) = +81.49 > reaction endothermic at 298 K